OpenStudy (anonymous):

Simplify the compound fractional expression. (Give your answer in factored form.)

OpenStudy (goformit100):

0k

OpenStudy (anonymous):

\[[\frac{ x }{ y }-\frac{ y }{ x}]/[\frac{ 1 }{ 5x^2 }-\frac{ 1 }{ 5y^2 }]\]

OpenStudy (anonymous):

i got 5x/-1 which is apparently wrong

OpenStudy (anonymous):

i just got the correct answer which is -5xy but thank you!

OpenStudy (goformit100):

keep it up :)